Segmented Reporting and Variances

Pittsburgh-Walsh Company (PWC) is a manufacturing company whose product line consists of lighting fixtures and electronic timing devices. The Lighting Fixtures Division assembles units for the upscale and mid-range markets. The Electronic Timing Devices Division manufactures instrument panels that allow electronic systems to be activated and deactivated at scheduled times for both efficiency and safety purposes. Both divisions operate out of the same manufacturing facilities and share production equipment.

   PWC’s budget for the year ending December 31, 20x1, follows and was prepared on a business segment basis under the following guidelines:

  1. Variable expenses are directly assigned to the incurring division.
  2. Fixed overhead expenses are directly assigned to the incurring division.
  3. The production plan is for 8,000 upscale fixtures, 22,000 mid-range fixtures, and 20,000 electronic timing devices. Production equals sales.

   PWC established a bonus plan for division management that required meeting the budget’s planned operating income by product line, with a bonus increment if the division exceeds the planned product line operating income by 10 percent or more.

PWC Budget
For the Year Ending December 31, 20x1
(In Thousands of Dollars)
Lighting Fixtures
Upscale Mid-Range Electronic Timing
Devices
Total
Sales $1,440 $ 770 $ 800 $ 3,010
Variable expenses:
Cost of goods sold (720) (439) (320) (1,479)
Selling and administrative (170) (60) (60) (290)
Contribution margin $ 550 $ 271 $ 420 $ 1,241
Fixed overhead expenses 140 80 80 300
Segment margin $ 410 $ 191 $ 340 $ 941

   Shortly before the year began, the CEO, Jack Parkow, suffered a heart attack and retired. After reviewing the 20x1 budget, the new CEO, Joe Kelly, decided to close the lighting fixtures mid-range product line by the end of the first quarter and use the available production capacity to grow the remaining two product lines. The marketing staff advised that electronic timing devices could grow by 40 percent with increased direct sales support. Increases above that level and increasing sales of upscale lighting fixtures would require expanded advertising expenditures to increase consumer awareness of PWC as an electronics and upscale lighting fixtures company. Kelly approved the increased sales support and advertising expenditures to achieve the revised plan. Kelly advised the divisions that for bonus purposes the original product-line operating income objectives must be met, but he did allow the Lighting Fixtures Division to combine the operating income objectives for both product lines for bonus purposes.

   Prior to the close of the fiscal year, the division controllers were furnished with preliminary actual data for review and adjustment, as appropriate. These preliminary year-end data reflect the revised units of production amounting to 12,000 upscale fixtures, 4,000 mid-range fixtures, and 30,000 electronic timing devices and are presented as follows:

PWC Preliminary Actuals
For the Year Ending December 31, 20x1
(In Thousands of Dollars)
Lighting Fixtures
Upscale Mid-Range Electronic Timing
Devices
Total
Sales $ 2,160 $140 $1,200 $ 3,500
Variable expenses:
Cost of goods sold (1,080) (80) (480) (1,640)
Selling and administrative (260) (11) (96) (367)
Contribution margin $ 820 $ 49 $ 624 $ 1,493
Fixed overhead expenses 140 14 80 234
Segment margin $ 680 $ 35 $ 544 $ 1,259

   The controller of the Lighting Fixtures Division, anticipating a similar bonus plan for 20x2, is contemplating deferring some revenues to the next year on the pretext that the sales are not yet final and accruing in the current year expenditures that will be applicable to the first quarter of 20x2. The corporation would meet its annual plan, and the division would exceed the 10 percent incremental bonus plateau in 20x1 despite the deferred revenues and accrued expenses contemplated.

Required:

1. Calculate the contribution margin, contribution margin volume, and sales mix variances. Enter your answers in dollars, rather than in thousands of dollars For example, enter "750,000" rather than "750". If required, round calculations to the nearest cent.

Contribution margin variance $
Contribution margin volume variance $
Sales mix variance $
Answer #1

As part 1 has already been answered, I have calculated the variances as required in Part 2.

Part 2)

The contribution margin variance is determined as below:

Contribution Margin Variance = Total Contribution Margin (Actuals) - Total Contribution Margin (Budget)

Substituting values in the above formula, we get,

Contribution Margin Variance = 1,493,000 - 1,241,000 = $252,000 (Favorable)

_____

The contribution margin volume variance is arrived as follows:

Contribution Margin Volume Variance = Budgeted Average Contribution Per Unit*(Actual Units Sold - Budgeted Unit Sales)

Here, Actual Units Sold = 12,000 + 4,000 + 16,000 = 46,000, Budgeted Unit Sales = 8,000 + 22,000 + 20,000 = 50,000 and Budgeted Average Contribution Per Unit = 1,241,000/50,000 = $24.82

Substituting these values in the above formula, we get,

Contribution Margin Volume Variance = 24.82*(46,000 - 50,000) = $99,280 (Unfavorable)

_____

The sales mix variance is calculated as below:

Sales Mix Variance = (Actual Units Sold for Upscale Fixtures - Budgeted Units Sold for Upscale Fixtures)*(Budgeted Contribution Per Unit for Upscale Fixtures - Budgeted Average Contribution Per Unit) +  (Actual Units Sold for Mi-Range Fixtures - Budgeted Units Sold for Mid-Range Fixtures)*(Budgeted Contribution Per Unit for Mid-Range Fixtures - Budgeted Average Contribution Per Unit) +  (Actual Units Sold for Electronic Timing Devices - Budgeted Units Sold for Electronic Timing Devices)*(Budgeted Contribution Per Unit for Electronic Timing Devices - Budgeted Average Contribution Per Unit)

Substituting values in the above formula, we get,

Sales Mix Variance = (12,000 - 8,000)*(550,000/8,000 - 24.82) + (4,000 - 22,000)*(271,000/22,000 - 24.82) + (30,000 - 20,000)*(420,000/20,000 - 24.82) = $362,520 (Favorable)

_____

Tabular Representation:

Contribution margin variance $252,000 Favorable
Contribution margin volume variance $99,280 Unfavorable
Sales mix variance $362,520 Favorable